# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ III-A2 Header based generation
Table I: Comparison of selected properties of packet and flow level data and their impact on data generation process
@@
Granularity High granularity offer more details of packet data e.g., packet header, payload. Low granularity as only a summary of the selected packet stream is given. Depending on the granularity requirement, the complexity of the generation task can vary. For example, packet levels with higher granularity may require complex model architectures to learn patterns with higher variability.
Amount of data High amount of data as each packet is considered Low amount of data as a summary of each flow is recorded [61] High-volume packet-level analysis captures detailed variations but may slow model convergence [61], while lower-variability flow-level data enables faster convergence with smaller datasets [89]
@@
Table II: Data collection tools at OSI layers of the network
OSI layer Tools used Basic details
Application Layer HTTP Proxies (e.g., Burp Suite [120], Fiddler [149]) Act as a proxy which captures or intercepts HTTP/HTTPS traffic. Different capturing modes are provided to capture local host or remote host traffic within the same local area network. [150]
Wireshark [159] Provides encrypted version of payload added at the application layer in hexadecimal notation.
Transport Layer Wireshark [159] Provides network packet captures related with various transport layer protocols (e.g., TCP, UDP, SCTP) with easy-to-handle Graphical User Interface (GUI). Can work in different operating systems.
Tcpdump [151] A lightweight Command Line Interface (CLI) tool for network packet capture that collects packets under the same protocols from the transport layer. Mostly supports Unix-like system.
Network Firewalls (Cisco ASA (Adaptive Security Appliance) [145, 146]) Sit between the internal network and external network (e.g., internet). Monitor and control packets particularly for detecting anomaly and malicious behavior.
Network Layer Router and Switch Monitoring (e.g., ManageEngine OpManager [95]) Especially designed for network layer device monitoring. Conduct port level analysis, device failures, and create switch/router level logs.
Wireshark and Tcpdump Both tools capture network layer protocols including IPv4, IPv6, ICMP, ARP etc.
Network IDS (e.g., Snort [139], Suricata [143]) Detect malicious packet streams with rule-based mechanisms while enabling packet capture, statistical analysis, and profiling.
Data Link Layer Wireshark and Tcpdump Both tools capture data link layer protocols including Ethernet, WiFi, etc.
Network TAPs (Test Access Points) [38, 142, 144] Physical device which can capture Layer 2 packet data. Copies the traffic through a given link to a monitoring device.
Physical Layer Network TAPs Can capture the raw electrical signal bit-by-bit before any further processing. Capturing happens in a non-intrusive manner without interfering with the signal flow
Software Defined Radios (SDR) (e.g.,RTL-SDR [112], HackRF One [53]) SDR contains its basic components as in software implementation replacing the traditional hardware components. This adds more flexibility in capturing and analyzing wireless signals from various technologies: WiFi, radio, Bluetooth, ZigBee, etc.
Packet and Flow level data: As shown in Fig. 2, header-based information is used to create mainly two data formats: Packet and Flow level data. In packet-level data, the raw network packet capture, usually in .pcap file format, is used for data generation [87, 45, 57, 64, 30]. As shown in Fig. 3, a given .pcap file is a time-series sequence of network packets and each record (or raw) of the packet capture contains multiple features from different layers described before in this section. Once a data generation model is trained with packet capture data, the model is expected to synthesize packet captures with the same properties and these properties can vary significantly based on user application, network bandwidth conditions, connection medium used [69, 87].
In flow level data, we create summary statistics of selected network packet stream (see Fig. 3). This packet stream, a.k.a. flow, is defined by a tuple of five elements: source and destination IP addresses, source and destination port addresses, and protocol [82]. In addition to these five elements, a given flow often contains summary statistics of the packet stream including the total number of packets, total bytes, average throughput, etc. in both uplink and downlink directions [35]. Generally, flow-level data generation models take a collection of such flow-level records as the input and synthesize similar records maintaining the same flow-level properties. Similar to packet-level data, flow-level properties vary based on similar conditions we highlighted in the previous paragraph.
Packet and flow-level data have their own advantages and disadvantages in terms of network traffic generation. We compare and contrast selected properties of these two data types and how those affect the data generation in Table I. For a better comparison, we assume packet-level and flow-level generated using the same packet capture.
III-A3 Signal generation
Data transmission signals are present in the Physical layer, where data is sent through either wired (e.g., Ethernet, fiber optics) or wireless (e.g., WiFi, radio waves, Bluetooth) channels. In these mediums, signals are transmitted as electrical or optic pulses and wireless waves. Compared to all other upper-layer data, generating the patterns in the physical layer is significantly challenging because of the high variability and no semantic meaning in the signals.
Despite these challenges, the generation of physical layer signal is attempted in [169, 111, 113, 134]. The authors in [111, 113] present a physical layer modulation and a coding scheme taking codeword indices or bits from radio signal as the input. They leverage conditional GANs for this purpose which compares two conditional probability distributions representing ground truth and approximated distributions. In [134], GANs are employed to generate adversarial signals capable of initiating spoofing attacks during wireless data transmission. The adversary intercepts these signals and constructs ground truth datasets by positioning a malicious receiver close to the legitimate receiver, enabling it to capture signals with nearly identical channel characteristics.
III-A4 Network usage conditions
As illustrated in Fig. 2, network usage conditions include various conditional attributes that influence network traffic transmission. For instance, at the Application layer, factors such as the type of application (e.g., web browsing or video streaming) or user behavior (e.g., whether usage occurs during the day or night) impact the overall nature of transmitted network traffic. For example, while video streaming patterns contain periodic chunks of bytes downloading, IoT traffic may show non-periodic pulses sending short data streams to the IoT servers [137]. At the Network and Transport layers, the available bandwidth plays a key role in determining network throughput. Therefore, generative models that incorporate these metadata elements can enhance the contextual relevance of synthetic data.
It is important to note that a given attribute may serve as a condition for one network traffic trace but not for another [87]. For example, the packet time (i.e., indicating the time of the day or peak and off-peak time) might influence available network bandwidth, thereby affecting synthetic traffic patterns as a conditional factor. However, this same information could also function as standard packet header information generated by the model which will be used in header-based data generation. Thus, conditional attributes are highly context-dependent and vary according to the data generation scenario.
Very little work incorporates explicit conditional network traffic generation. In [87], conditional data (i.e., introduced as metadata by the authors) is collected as a ground truth component alongside related network traffic data. For example, in one of their datasets, which includes network packet-level features such as round-trip times and packet loss rates, factors like internet connection technology, ISP, and client location are considered metadata. They utilize an auxiliary GAN model to generate this metadata, which is then used as conditional input to generate the main network traffic data.
III-B Data collection devices/vantage points
Collecting a sufficient amount of network traffic data is important to have high-quality synthetic data. This can be achieved by deploying appropriate data collection tools at the required network layer. Table II reports a basic analysis of such data collection tools at different network layers. Note that several tools such as Wireshark, Tcpdump, and network TAPs provide network data information at different network OSI layers. For example, Wireshark provides packet content from the Application layer (encrypted payload) to the Data link layer. While we have found HTTP proxy tools that can collect HTTP/HTTPS, to the best of our knowledge no related work could be found in HTTP data generation.
IV Statistical methods based generation
@@
First, captured network traffic samples are processed to extract statistical properties of traffic, such as arrival rates, inter-arrival times, packet sizes, etc., and based on the characteristics of the extracted features, a suitable distribution/model (such as Poisson process, Markov Model, etc. that are discussed in this section) is chosen. Next, a model/distribution will be generated by estimating the corresponding parameters to fit the captured data, and the generated distribution/model will be validated against the real traffic. Once a suitable model is established, it can be used to predict future traffic patterns by changing the model parameters as necessary. These models were used to predict future traffic behavior and inform network planning and performance analysis. Additionally, synthetic traffic traces could be generated based on these models to simulate various scenarios and evaluate network performance.
@@
Table III: Summary of works leveraging statistical methods.
Work Application scenario Method(s) used Data Artifacts
OSI layer Format/Type available
Paxson et al. [117] Modeling wide area traffic Poisson Model Transport Session/ Flow/ Packet No
Ammar et al. [7] Generating realistic IP traffic Poisson Pareto Burst Process Network Flow Yes
Yoonsen et al. [172] Generating data center network traffic Poisson Short-Noise Model Network Flow No
Arfeen et al. [11] Modeling traffic in Internet access and backbone core networks Weibull Distribution Transport/ Network Session/ Flow/ Packet No
Lee et al. [79] Modeling TCP/IP traffic of wireless local area networks Piecewise Weibull Distribution with Canonical Multinomial Cascade Model Transport/ Network Flow/Packet No
Arfeen et al. [12] Modeling Internet traffic Weibull Distribution (two-parameter) Transport/ Network Session/ Flow/ Packet No
Muscariello et al. [105] Modelling Internet traffic 5-parameter Hierarchical Markov Modulated Poisson Process Transport/ Network Session/ Flow/ Packet No
Xie et al. [163] Modeling time-varying oscillation behavior of Internet traffic Nested Hidden Markov Models with Variable State-duration Transport/ Network Arrival rate process No
Dainotti et al. [43] Source-based modeling of internet traffic st packet level Hidden Markov models Application Packet No
Garcia et al. [50] Detecting malicious network behavior Markov Chains Transport Flow No
Shu et al. [135] Predict network traffic FARIMA Network Packet No
Zhou et al. [179] Predicting network traffic ARIMA/GARCH Transport Packet No
Pashko et al. [116] Modeling of self-similar traffic and loading process of telecommunication networks Fractional Brownian Motion Not specified Not specified No
Meng et al. [98] Modeling control plane traffic of 5G networks Two-level Hierarchical State-machine Application Control plane events No
Cieslak et al. [32] Handling imbalance in IDS K-nearest Neighbors Not specified Not specified No
Vu et al. [153] Handling data imbalance for encrypted traffic identification K-nearest Neighbors Network Flow No
IV-A Poisson process
@@
Early research in modeling network traffic and synthesizing focused on using the Poisson process to accurately model the dynamics of network traffic. A major factor for utilizing the Poisson process, rooted in the principles of statistical independence, proved effective for certain scenarios [85]. Research was done to explore the limitations of Poisson modeling and identified scenarios where it failed to accurately represent the observed traffic patterns [117]. In the early 1990s, two groundbreaking papers revealed that both Wide Area Network (WAN) and Local Area Network (LAN) traffic exhibit Long Range Dependence (LRD) and self-similarity [80], [117]. In addition, the authors of [117] show that Poisson processes are valid only for modeling the arrival of user sessions (TIELWET connections, FTP control connections), but they perform poorly for other WAN arrival processes. Therefore, it became evident that the Poisson process was inadequate for this task. Recognizing the limitations of the Poisson process, researchers initially explored modified approaches based on the Poisson process to capture LRD and self-similarity properties.
@@
One such alternative model used is the Poison Pareto Burst Process (PPBP). Ammar et al. [7] introduced a new tool for generating realistic internet traffic named NS-3, an open-source discrete-event network simulator for internet systems. By utilizing the Poisson Pareto Burst Process, the paper focused on capturing the LRD characteristics observed in real-world traffic. The traffic model was validated using an autocorrelation function plot that revealed the LRD property of real-world internet traffic.
@@
The limitations of the Poisson process were more evident when modeling high-speed network traffic, particularly real-time data traffic modeling for next-generation networks. For example, Liji et al. [85] demonstrated that the Stationary Poison Increment Process can only model Short Range Dependence (SRD) but not LRD. To address this limitation, the authors proposed using second-order self-similarity models, such as fractional Gaussian noise and fractional ARIMA processes, as a more appropriate approach.
In the meantime, researchers also explored modeling data center network traffic using poisson processes. To better simulate realistic traffic in data center environments, the generation of flow-level network traffic matrices based on the poisson shot-noise model is proposed in [172]. By incorporating factors such as flow arrival rates, intra-rack traffic ratios, flow sizes and durations, the poisson shot-noise process offers a more accurate representation of traffic patterns in data centers.
IV-B Weibull distribution

### v2

@@ II-B Transport, Network, and Data Link layers (L4–L2): Packet headers
Table II: Comparison of selected properties of packet and flow level data and their impact on data generation process
@@
Granularity High granularity offer more details of packet data e.g., packet header, payload. Low granularity as only a summary of the packet stream is given. Generation complexity depends on granularity; packet-level data with higher variability often requires more complex models.
Amount of data High amount of data as each packet is considered Low amount of data as a summary of each flow is recorded [71] High-volume packet-level analysis captures detailed variations but may slow model convergence [71], while lower-variability flow-level data has faster convergence with smaller datasets [104]
@@
The most widely used data representations in network traffic generation are derived from packet headers at the Transport, Network, and Data Link layers. This is largely because header fields are observable at many network vantage points, including routers, switches, firewalls, and monitoring appliances, even when payloads are encrypted. As a result, header-level data offers a practical and broadly accessible view of network behavior across operational environments.
Header information captures concise but expressive summaries of packet properties, such as packet size, protocol type, source and destination addresses, port numbers, and timing information. These features are sufficient for many downstream tasks, particularly in machine learning–driven traffic analysis, including traffic classification, anomaly detection, and performance modeling [110, 76, 43]. Consequently, most traffic generation studies operate on header-derived data rather than full payloads or physical-layer signals.
While headers at different layers expose distinct fields, the most commonly used attributes for data generation include:
  1. 1.
    Transport layer (L4): Transmission Control Protocol (TCP) and User Datagram Protocol (UDP) headers provide source and destination port numbers, control flags (e.g., SYN, ACK, FIN, RST), and segment-level information, which are widely used to characterize connection behavior and session dynamics [76, 197].
  2. 2.
    Network layer (L3): The Internet Protocol (IP) header exposes packet length, source and destination IP addresses, time-to-live (TTL), and upper-layer protocol identifiers, enabling the modeling of routing behavior and end-to-end communication patterns [197, 190].
  3. 3.
    Data link layer (L2): Medium Access Control (MAC) addresses, EtherType values, and frame length provide visibility into local network topology, link-layer protocol usage, and broadcast or multicast behavior [44].
Some studies focus on data extracted from a single layer, such as frame length at the Data Link layer [157, 156], while others combine fields across multiple layers to construct multivariate representations. For example, in [197, 76], authors extract source and destination IP addresses and protocol identifiers from the Network layer together with port information from the Transport layer, enabling richer representations of communication patterns.
As shown in Fig. 3, header-based information is primarily organized into two data representations: packet-level and flow-level data, which differ in aggregation granularity but are derived from the same observable header fields.
II-B1 Packet-level representation
Packet-level data consists of raw packet captures, typically stored in .pcap format [102, 51, 68, 76, 33]. Each record in a packet trace corresponds to an individual packet and contains header fields from multiple layers, along with optional payload data (Fig. 4). From a visibility standpoint, packet-level representations preserve fine-grained temporal dynamics and per-packet variations, making them suitable for modeling detailed protocol behavior, timing patterns, and burstiness. However, they also result in large data volumes and higher processing overhead.
II-B2 Flow-level representation
Flow-level data aggregates packets into logical communication units, or flows, typically defined by a five-tuple consisting of source IP, destination IP, source port, destination port, and protocol [96]. In addition to this identifying tuple, flows often include summary statistics such as total packet count, total bytes, duration, and average throughput in both directions [38]. This representation is widely standardized through telemetry frameworks such as Cisco NetFlow [39], sFlow [132], and the IETF IP Flow Information Export (IPFIX) protocol [40], enabling interoperable and vendor-agnostic traffic reporting across large network infrastructures. Beyond simple aggregation, flow records support the construction of statistical fingerprints of network behavior—capturing distributions of packet sizes, inter-arrival times, directional asymmetries, and temporal activity patterns—which are frequently leveraged in anomaly detection, intrusion detection, and traffic classification systems. From a visibility perspective, flow records therefore provide a compact yet analytically expressive abstraction of traffic behavior, making them suitable for large-scale monitoring, long-term analysis, and scalable traffic generation, albeit at the cost of losing fine-grained packet-level dynamics.
Packet-level and flow-level representations exhibit complementary strengths and limitations in terms of data volume, granularity, processing requirements, and noise sensitivity. Table II compares selected properties of these two data types and highlights their implications for traffic generation. For consistency, we assume that both representations are derived from the same underlying packet capture.
Data collection mechanisms: A wide range of tools are used to collect packet header and flow data across different layers of the network stack. Packet capture tools such as Wireshark [181] and Tcpdump [167] provide multi-layer visibility by extracting header information, including IP addresses at the Network layer and port details at the Transport layer. At the Transport layer, network firewalls (e.g., Cisco ASA) are commonly deployed to monitor, filter, and log connection-level activity. At the Network layer, router and switch monitoring platforms such as ManageEngine OpManager [111] are widely used to collect traffic statistics, perform port-level analysis, and detect device failures. At the Data Link layer, hardware-based network Test Access Points (TAPs) [162] and switch-based port mirroring techniques are employed to duplicate traffic from specific links or ports, enabling passive header-level observation without interfering with traffic flow.
II-C Physical layer (L1): Signal-level representations
At the Physical layer, network traffic is represented as electrical, optical, or radio-frequency (RF) signals transmitted over wired or wireless media. Unlike higher-layer representations, physical-layer data consists of waveform-level measurements that encode information through modulation, amplitude, phase, and frequency variations. These signals do not directly expose packet headers or application semantics, as higher-layer protocol structures are reconstructed only after signal demodulation and decoding. Furthermore, physical-layer signals exhibit high variability due to channel effects such as noise, interference, attenuation, and multipath propagation.
From a data visibility perspective, physical-layer representations are accessible only at specialized observation points, such as hardware taps on wired links or radio receivers in wireless environments. As a result, the collected data typically consists of raw time-domain signal samples, frequency-domain spectral representations, or derived measurements reflecting transmission and channel characteristics.
Several prior works have explored synthetic generation of such signal-level representations across different wireless technologies [192, 35, 54, 150, 154, 180]. These studies operate directly on RF waveform data or their spectral representations to reproduce signal characteristics observed in Wi-Fi, radar, cellular, and RFID systems. In addition, higher-frequency signals such as millimeter-wave (mmWave) and Terahertz communications have been considered in [16]. Beyond waveform generation, related physical-layer attributes such as Channel State Information (CSI), which characterizes signal propagation and channel effects, have also been synthetically generated to emulate realistic transmission environments [173]. Furthermore, works such as [128, 129] consider generation at the modulation and coding level using physical-layer bit sequences as input representations.
Data collection devices: At the physical layer, raw signal data is collected using specialized hardware capable of capturing transmissions prior to higher-layer decoding. In wired networks, Network Test Access Points (TAPs) provide non-intrusive access to electrical or optical signals directly from communication links. In wireless environments, Software Defined Radios (SDRs) such as RTL-SDR and HackRF [131, 63] serve as primary data collection tools, enabling flexible capture of RF signals across a wide range of wireless technologies, including Wi-Fi, Bluetooth, ZigBee, and cellular systems. Additionally, wireless interfaces operating in monitor mode, supported by tools such as Aircrack-ng [163], enable the collection of low-level transmission data and signal metadata associated with wireless communication.
II-D Contextual and conditional attributes across layers
Certain factors influencing network traffic, such as user behavior, application type, available bandwidth, time of day, and connection medium, are not directly observable at any single OSI layer. Instead, these attributes reflect external context or operational conditions that shape traffic patterns across multiple layers. For example, the application in use (e.g., video streaming versus web browsing) is defined at the Application layer but manifests indirectly through packet sizes, flow durations, and throughput at lower layers. Similarly, the connection medium (e.g., fiber optics versus wireless) constrains available bandwidth and latency, influencing observable characteristics at the Transport and Network layers.
From a data visibility perspective, these attributes are typically not extracted directly from packet captures or flow records. Instead, they are inferred, logged separately by network management systems, or known a priori in controlled environments. As such, they are more appropriately treated as conditional attributes or meta-data that guide traffic generation rather than as primary data representations.
The synthetic generation of such contextual attributes—including application selection, user activity patterns, access technology, and usage timing—is essential for enabling realistic, controlled, and scalable traffic generation frameworks. Prior work [102, 197] has demonstrated the feasibility of generating limited sets of such meta-data and using them as conditional inputs to traffic synthesis pipelines. However, these efforts are restricted to narrow attribute spaces. More broadly, systematic generation and integration of richer, multi-dimensional contextual attributes remain largely unexplored, representing an open challenge for controlled and policy-aware traffic generation.
In this section, we presented the different types of network traffic data across the protocol stack, ranging from physical-layer signal representations to higher-layer packet and flow-level abstractions. Each data type provides a distinct level of visibility into network activity, with varying structural characteristics, semantic richness, and collection mechanisms. These differences directly influence how network traffic can be modeled and synthesized. Building on this foundation, the following sections examine synthetic data generation methods, focusing on the techniques and approaches used to generate realistic network traffic data across these different representation levels.
III Statistical methods based generation
@@
First, captured network traffic samples are processed to extract statistical properties of traffic, such as arrival rates, inter-arrival times, packet sizes, etc., and based on the characteristics of the extracted features, a suitable distribution/model (such as PP, MM, etc. that are discussed in this section) is chosen. Next, a model/distribution will be generated by estimating the corresponding parameters to fit the captured data, and the generated distribution/model will be validated against the real traffic. Once a suitable model is established, it can be used to predict future traffic patterns by changing the model parameters as necessary. These models were used to predict future traffic behavior and inform network planning and performance analysis. Additionally, synthetic traffic traces could be generated based on these models to simulate various scenarios and evaluate network performance.
@@
Table III: Summary of works leveraging statistical methods with pros and cons.
Work Application scenario Method(s) used Data Pros Cons
OSI layer Format/Type
[136] Wide area traffic $P$ Transport Session/Flow/Packet Simple and tractable Misses LRD
[7] IP traffic PPBP Network Flow Captures LRD Assumes fixed burst pattern
[198] Data center traffic generation PSN Network Flow Models realistic DCN flows Less suited for non-DCN traffic
[195] D2D mmWave modeling Poisson Cluster Process Physical mmWave Captures spatial clustering Assumes idealized distribution
[98] Downlink D-MIMO modeling Poisson point process Physical mmWave Models RAU/UE randomness Ignores temporal patterns
[11] Access/core Internet traffic WD Transport/ Network Session/ Flow/ Packet Fits heavy/light tails Misses correlation structures
[93] WLAN TCP/IP modeling Piecewise Weibull + Cascade Transport/ Network Flow/Packet Captures flow inter-arrivals Can’t model intra-flow arrivals
[12] Internet traffic modeling WD Transport/ Network Session/ Flow/ Packet Links tiers via parameters Needs tier-specific tuning
[137] Indoor positioning WD Physical RSSI Fits RSSI well Needs precise parameters
[121] Internet traffic modeling 5-param Hierarchical MMPP Transport/ Network Session/ Flow/ Packet Captures key traffic metrics Lacks full LRD modeling
[185] Oscillatory traffic modeling Nested HMM (var. duration) Transport/ Network Arrival rate process Matches oscillations Higher complexity
[47] Packet-level modeling HMMs Application Packet Predicts from little data Short-term only
[59] Malicious behavior detection Markov Chains Transport Flow Good for botnet C&C Botnet-specific behavior
[79] Wireless link simulation HMM + MMB Physical Wireless Captures multi-scale dynamics Needs rich measurements
[155] Traffic prediction ARFIMA Network Packet Models SRD+LRD Fitting is complex
[206] Traffic prediction ARIMA/GARCH Transport Packet Captures SRD, LRD, multifractal Many parameters
[135] Self-similar traffic modeling FBM Not specified Not specified Natural self-similarity Weak short-term modeling
[114] 5G control-plane modeling HSM Application Control plane events Captures bursty deps Complex to deploy
[36] IDS imbalance handling KNN Not specified Not specified Aids minority detection Risk of overfitting
[170] Encrypted traffic classification KNN Network Flow Combines sampling methods Dependent on sampling quality
III-A Poisson process (PP)
@@
Early research in modeling network traffic and synthesizing focused on using the PP to accurately model the dynamics of network traffic. A major factor for utilizing the PP, rooted in the principles of statistical independence, proved effective for certain scenarios [100]. Research was done to explore the limitations of Poisson modeling and identified scenarios where it failed to accurately represent the observed traffic patterns [136]. In the early 1990s, two groundbreaking papers revealed that both WAN and LAN traffic exhibit Long Range Dependence (LRD) and self-similarity [94], [136]. In addition, the authors of [136] show that PPs are valid only for modeling the arrival of user sessions (Telecommunication Network [TELNET] connections, File Transfer Protocol [FTP] control connections), but they perform poorly for other WAN arrival processes. Therefore, it became evident that the PP was inadequate for this task. Recognizing the limitations of the PP, researchers initially explored modified approaches based on the PP to capture LRD and self-similarity properties.
@@
One such alternative model used is the Poisson Pareto Burst Process (PPBP). Ammar et al. [7] introduced a new tool for generating realistic internet traffic named NS-3, an open-source discrete-event network simulator for internet systems. By utilizing PPBP, the paper focused on capturing the LRD characteristics observed in real-world traffic. The traffic model was validated using an autocorrelation function plot that revealed the LRD property of real-world internet traffic.
@@
The limitations of the PP were more evident when modeling high-speed network traffic, particularly real-time data traffic modeling for next-generation networks. For example, Liji et al. [100] demonstrated that the Stationary Poison Increment Process can only model Short Range Dependence (SRD) but not LRD. To address this limitation, the authors proposed using second-order self-similarity models, such as fractional Gaussian noise and Autoregressive Fractionally Integrated Moving Average (ARFIMA) processes, as a more appropriate approach. In the meantime, researchers also explored modeling data center network traffic using PPs. To simulate traffic in data center environments, the generation of flow-level network traffic matrices based on the Poisson Shot-Noise (PSN) model is proposed in [198]. By incorporating factors such as flow arrival rates, intra-rack traffic ratios, flow sizes and durations, the PSN process offers a more accurate representation of traffic patterns in data centers.
Researchers have investigated modeling signals or metrics of physical layer. In [195], the authors study the performance of mmWave communications in clustered device-to-device (D2D) networks, where D2D transceivers are spatially modeled using a Poisson Cluster Process. Each cluster features multi-antenna devices, and mmWave links are used by active D2D transmitters to serve nearby receivers. The paper proposes and analyzes three user association strategies: uniformly distributed, nearest, and closest line-of-sight (LOS) D2D transmitter models. In [98], the authors analyze a downlink distributed millimeter wave massive Multiple-Input Multiple-Output (MIMO [D-MIMO]) system where radio access units (RAUs) and user equipments (UEs) are modeled as Poisson point processes. A hybrid precoding scheme based on antenna array response vectors is applied, and the impact of the RAU-to-UE ratio on system performance is examined. Two RAU allocation strategies—Distance-based and Signal-to-Interference-plus-Noise Ratio (SINR)-based—are proposed, and a lower bound on the asymptotic average spectral efficiency is derived for the Distance-based approach.
Limitations: Taken together, the studies discussed in this subsection indicate that the core limitation of Poisson-based traffic models lies in their structural assumptions. The independence and stationary increment properties inherent to PPs conflict with empirical observations of long-range dependence and self-similarity reported for WAN and LAN traffic [94, 136]. While PP-derived extensions such as PPBP and PSN partially mitigate these issues by introducing burst-level dynamics or flow aggregation [7, 198], they still rely on Poisson arrivals at a fundamental level. Consequently, these models remain constrained in their ability to represent correlated behavior across multiple time scales, particularly outside the specific scenarios for which they are designed.
Summary: In summary, PPs form the conceptual baseline for statistical network traffic generation and have played a crucial role in shaping early modeling efforts. Their simplicity enables clear analytical insights and supports extensions into spatial and flow-level modeling at higher layers [195, 98]. However, the empirical evidence presented in this subsection consistently shows that realistic traffic behavior requires mechanisms beyond memoryless arrivals. As a result, PP-based models are best viewed as building blocks or components within more expressive frameworks rather than complete solutions for modern traffic synthesis.
III-B Weibull distribution (WD)